OpenStudy (captbrazil):

Find the derivative of f(x) = -7/x at x = -3.

OpenStudy (captbrazil):

I got 7/3 as the answer but it said its wrong

OpenStudy (mathmale):

It may help you to re-write f(x) = -7/x as \[f(x)=-7x ^{-1}.\]

OpenStudy (captbrazil):

ok so, 1/-7x?

OpenStudy (mathmale):

Focus on \[x ^{-1}\]. What's the derivative of \[x ^{-1}?\]

OpenStudy (mathmale):

Hint: Use the Power Rule for Differentation.

OpenStudy (mathmale):

\[\frac{ d }{ dx }x^n=nx ^{n-1}\]

OpenStudy (mathmale):

So, \[\frac{ d }{ dx }x ^{-1}=?\]

OpenStudy (captbrazil):

never saw that before, online class skipped over it completely Let me think

OpenStudy (mathmale):

What is n in\[\frac{ d }{ dx }x ^{-1}~?\]

OpenStudy (captbrazil):

-1

OpenStudy (captbrazil):

so we get -1x^-2?

OpenStudy (mathmale):

Again, this is the Power Rule for Differentiation. Yes, n=-1. Now, please find\[\frac{ d }{ dx }x ^{-1}\]

OpenStudy (mathmale):

Yes, you have that correct.

OpenStudy (mathmale):

But remember that your original problem began with "-7." You must keep that "-7". So, what is your final result?

OpenStudy (captbrazil):

-7x^-2?

OpenStudy (mathmale):

\[\frac{ d }{ dx }[-7x ^{-1}]=?\]

OpenStudy (mathmale):

You have -7 as your coefficient, and your n is = -1. What is (-7)(-1)?

OpenStudy (captbrazil):

oops you're right, its +7x

OpenStudy (mathmale):

Your +7 is correct, but your "x" is not.

OpenStudy (mathmale):

\[\frac{ d }{ dx }[-7x ^{-1}]=?\]

OpenStudy (mathmale):

You must use the "Power Rule for Differentiation."

OpenStudy (captbrazil):

+7x^-2

OpenStudy (mathmale):

Yes, basically that's correct. I'd prefer that you write 7x^(-2), so that it is clear to the reader that your exponent is negative. Now, let x=-3 in your final result.

OpenStudy (mathmale):

Be certain to enclose your -3 inside parentheses: 7(-3)^(-2)=?

OpenStudy (mathmale):

Note that you MUST do the exponentiation first here, and only then do the multiplication.

OpenStudy (captbrazil):

7*(-3)^(-2) = 7 * 1/9 = 7/9

OpenStudy (mathmale):

beautiful.

OpenStudy (captbrazil):

thank you very much

OpenStudy (mathmale):

my pleasure!